Solve each quadratic equation by factoring and applying the 0 product principle.
2x^2-x=1
answers= -5,2
3x^2-7x=0
answers= 0,7/3
Could someone please write out the steps for me?
Thank you!

Answer by Nate(3500)

2x^2-x=1 original equation
2x^2-x-1=0 set the equation to zero
2x^2-2x+x-1=0 the product leading coefficient and the constant is -2 (two numbers that add up to -1 and are multiplied to get the -2 is: -2 and 1)
(2x^2-2x)+(x-1)=0 start to combine
2x(x-1)+1(x-1)=0 take out the factors
(2x+1)(x-1)=0 set them to zero
2x+1 = 0 ; 2x = -1 ; x = -1/2
x-1 = 0 ; x = 1
1 and (-1/2) are your answers
-------------------------------------------------------------------------------
3x^2-7x=0
x(3x-7)=0 set them to zero
x=0
3x-7=0 ; 3x=7 ; x=7/3
0 and (7/3) are the answers